OpenStudy (anonymous):

1/2 - 6 = 14 r = 40 r = 34 r = 20 r = 16

OpenStudy (anonymous):

@Nurali

OpenStudy (jhannybean):

^ thats what i'm wondering. unless the person forgot to put it in the proper form?

OpenStudy (anonymous):

sorry I ment to put this lol r/2 - 6 = 14*

OpenStudy (nurali):

r/2 - 6 = 14 r/2 =14+6 r/2 =20 r=20*2 r=40

OpenStudy (anonymous):

thanks! can you help me on more?

OpenStudy (anonymous):

multiply 2 at both side

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

k thanks im gonna just tag you in it